OpenStudy (anonymous):

Help me find the solution to 64^(x+3) = 42^(2x)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (2^6)^{x + 3} = (2^2)^{2x} }\) \(\Large \color{purple}{\rightarrow 2^{6x + 18} = 2^{4x} }\) Do you know how to solve further?

Parth (parthkohli):

Just create an equation of the exponents when the bases are same. \(\Large \color{purple}{\rightarrow 6x + 18 = 4x }\) \(\Large \color{purple}{\rightarrow 18 = -2x }\) \(\Large \color{purple}{\rightarrow x = -9 }\)